OpenStudy (anonymous):

A bag contains 13 dark chocolates, 16 white chocolates, and 11 milk chocolates. Maggie’s class teacher, Jill, is very strict, and the probability that she offers her students chocolates is 0.4. If Jill offers Maggie a chocolate from the bag, what is the probability that she randomly picks either a white chocolate or a milk chocolate?

OpenStudy (abb0t):

@dan815

OpenStudy (anonymous):

13/0.4=32.5 for dark chocolates. 16/0.4=40 for white chocolates. 11/0.4=27.5 for milk chocolates. So , since the white chocolates are mostly what there is in the bag , therefore its more probable for you too pick the one that carries more, so therefore its 40% of whats in the bag . So its white chocolate.

OpenStudy (anonymous):

the choices are 11/10 27/40 4/7 4/10 or 13/40 @lalaxxo so i'm a little confused as to which answer it would be

OpenStudy (anonymous):

13/40

OpenStudy (anonymous):

its 13/40 since those are the 2 most probable to pick .

OpenStudy (anonymous):

thank you!!